{F} The equation for a displacement ๐‘ (๐‘š), at a time ๐‘ก(๐‘ ) by an object starting at a displacement of ๐‘ 0 (๐‘š), with an initial velocity ๐‘ข(๐‘š๐‘  โˆ’1 ) and uniform acceleration ๐‘Ž(๐‘š๐‘  โˆ’2 ) is: ๐‘  = ๐‘ 0 + ๐‘ข๐‘ก + 1 2 ๐‘Ž๐‘ก 2 A projectile is launched from a cliff with ๐‘ 0 = 30 ๐‘š, ๐‘ข = 55 ๐‘š๐‘  โˆ’1 and ๐‘Ž = โˆ’10 ๐‘š๐‘  โˆ’2 . The tasks are to: a) Plot a graph of distance (๐‘ ) vs time (๐‘ก) for the first 10s of motion. b) Determine the gradient of the graph at ๐‘ก = 2๐‘  and ๐‘ก = 6๐‘ . c) Differentiate the equation to find the functions for: i) Velocity (๐‘ฃ = ๐‘‘๐‘  ๐‘‘๐‘ก) ii) Acceleration (๐‘Ž = ๐‘‘๐‘ฃ ๐‘‘๐‘ก = ๐‘‘ 2 ๐‘  ๐‘‘๐‘ก2 ) d) Use your results from part c to calculate the velocity at ๐‘ก = 2๐‘  and ๐‘ก = 6๐‘ . e) Compare your results for part b) and part d). f) Find the turning point of the equation for the displacement ๐‘  and using the second derivative verify whether it is a maximum, minimum or point of inflection. g) Compare your results from f) with the graph you produced in a).


1
Expert's answer
2022-02-14T16:53:40-0500
"s(t)=30+55t-\dfrac{10t^2}{2}, m"

a)


"s(t)=30+55t-5t^2, 0\le t\le10"


b)


"grad \ s=\dfrac{s_2-s_1}{t_2-t_1}"

"t=2 , s(2)=120"


"t=0 , s=50"


"grad \ s|_{t=2}=\dfrac{120-50}{2-0}=35(m/s)"

"t=6 , s(6)=180"


"t=0 , s=210"

"grad \ s|_{t=6}=\dfrac{180-210}{6-0}=-5(m/s)"


c)

i)


"v(t)=\dfrac{ds}{dt}=55-10t, m/s"

ii)


"a(t)=\dfrac{dv}{dt}=\dfrac{d^2s}{dt^2}=-10\ m/s^2"

d)


"v(2)=55-10(2)=35(m\/s)"

"v(6)=55-10(6)=-5(m\/s)"

e) The results are the same.


f) The turning point of the equation for the displacement

"\dfrac{ds}{dt}=0=>55-10t=0"

"t=5.5 s"

"s(5.5)=30+55(5.5)-5(5.5)^2=181.25(m)"

Turning point is "(5.5, 181.25)"


"\dfrac{d^2s}{dt^2}=-10<0"

Point "(5.5, 181.25)" is a maximum.


g) The results are the same.